3.18. Find the mean and variance of the given PMF pr)-1/k, where - 1,2,3,, k.
Answer #1

here as we know that mean E(X)=\sumxP(x)=1*(1/k)+2*(1/k)+3*(1/k)+4*(1/k)+...k*(1/k)

E(X)=(1/k)*(1+2+3+4+...k)=(1/k)*(k(k+1))/2

E(X)=(k+1)/2

E(X2)=\sumx2P(x)=(1/k)*(12+22+32+..+k2)=(1/k)*(k(k+1)(2k+1))/6 =(k+1)*(2k+1)/6

therefore Variance =E(X2)-(E(X))2 =(k+1)*(2k+1)/6-(k+1)2/4